OpenStudy (anonymous):

At a town fair, 9 tickets cost twelve dollars. How many tickets could you expect to receive for eight dollars, assuming there is no discount for buying more tickets? 4 5 6 7

OpenStudy (texaschic101):

9/12 = x/8 (9 tickets to $12 = x tickets to $8) now cross multiply... (12)(x) = (9)(8) 12x = 72 Now divide both sides by 12, leaving you with just x on the left side. Don't forget that what you do on one side of an equation, you have to do on the other side. Can you solve for x now ? (x being the number of tickets $8 will buy)